If a county's labor is paid a total of 6 billion, it's capital is paid a total of 2 billion, and profits are zero, what is the level of output?
1
Expert's answer
2019-12-25T12:12:29-0500

If a county's labor is paid a total of 6 billion, its capital is paid a total of 2 billion, and profits are zero, then its total cost is TC = 6 + 2 = 8 billion, total revenue is TR = TC = 8 billion. The level of output is Q = TR/P = 8/P, so we need to know price P to find the level of output Q.